TheSmartOne (thesmartone):

Consider the function \( h(x) = \frac{1}{4}x^4 - \frac{5}{3}x^3 + 3x^2 + 4x\).Find all values of x where the tangent line to the graph of y = h(x) is parallel to the line y = 4x + 3.

TheSmartOne (thesmartone):

Consider the function \( h(x) = \frac{1}{4}x^4 - \frac{5}{3}x^3 + 3x^2 + 4x\).Find all values of x where the tangent line to the graph of y = h(x) is parallel to the line y = 4x + 3.

TheSmartOne (thesmartone):

So y = 4x + 3 has a slope of 4 So h'(x) = 4

TheSmartOne (thesmartone):

h'(x) = x^3 - 5x^2 + 6x + 4 and since h'(x) = 4 x^3 - 5x^2 + 6x + 4 = 4 x^3 - 5x^2 + 6x = 0 x(x^2 - 5x + 6) = 0 x = 0, 2, 3

OpenStudy (sshayer):

\[\frac{ dh }{ dx }=4*\frac{ 1 }{ 4 }x^3-3*\frac{ 5 }{ 3 }x^2+6x+4=x^3-5x^2+6x+4\] you are correct find x

satellite73 (satellite73):

fortunately for you that is easy to solve, although it is usually hard to solve a cubic equatoin

TheSmartOne (thesmartone):

Yeah, thanks all c: